You plan to retire with $8,000,000 in 38 years. How much should you deposit each month...

  1. You plan to retire with $8,000,000 in 38 years. How much should you deposit each month into an account that pays 9% annual rate compounded monthly? (Note: Compounding frequency is the same as the deposit frequency. Answers are rounded.)

    $2,056

    $1,080

    $1,198

    $1,124

    $1,276

    $1,709
Answer #1

Future value=$8000000
Present value=$0
Monthly interest rate=(9%)/2=0.045
Time period=38 years
When the interest rate is compounded monthly, the number of periods=38*12=456
We can calculate the monthly payment or deposit using excel.

8000000 1 Future value 2 Present value 3 Number of periods 4 Interest rate 5 Payments 456 0.0075 ($2,056.08)

1 Future value 8000000 2 Present value 3 Number of periods =38*12 4 Interest rate =9%/12 Payment: =PMT(Q4,Q3, Q2,01)

Hence, the monthly deposit is $2056 (Rounded to the nearest whole number)